# Functional Regression with Nonstationarity and Error Contamination: Application to the Economic Impact of Climate Change

### v1

@@ 5.2 Empirical Applications: Economic Impact of Climate Change
This section presents an empirical application on the economic impact of climate change using our proposed method. We consider appropriate transformations of the probability densities of gross regional product (GRP) growth rates ($y_{t}$), as a measure of regional economic activity, and land temperature anomalies ($x_{t}$), commonly used as indicators of climate change in the literature. These time series are expected to be nonstationary, contaminated by measurement errors, and to exhibit nonzero unconditional means. Accordingly, we apply model (24), with $u_{t}$ absorbing measurement error in the dependent variable.
Extensive evidence in the climate-economics literature shows that climate resilience depends on a country’s wealth and reliance on climate-sensitive industries such as agriculture and manufacturing (e.g., Dell et al.,, 2012; Burke et al.,, 2015; Newell et al.,, 2021; Cruz and Rossi-Hansberg,, 2023). Wealthier countries adapt more effectively to harsh environmental conditions, highlighting the spatial heterogeneity of climate-change impacts. Using our model, we investigate statistical evidence supporting this relationship, along with the general economic impact of climate change.
@@ 5.2.1 Raw data and functional data in analysis
We use non-infilled gridded land temperature anomaly data from a collaborative product of the Climatic Research Unit at the University of East Anglia, the Met Office Hadley Centre, and the National Centre for Atmospheric Science (CRUTEM.5.0.2.0, Osborn et al.,, 2021). We estimate spatially distributed temperature anomaly densities for 1951–2019 using a Gaussian kernel with Silverman’s bandwidth. To avoid COVID-19–related distortions in $y_{t}$, data from 2020 onward are excluded. At each time $t$, the distribution’s support is restricted to the range containing 99% of the total probability mass, [5.80,6.68][-5.80,6.68], thereby excluding outliers as in Chang et al., (2020).
For the GRP growth rates, we employ both the real GRP data of Wenz et al., (2023) for the period 1960–2019 and the real GDP in millions of 2021 international dollars, converted using Purchasing Power Parities, from the Conference Board Total Economy Database (TED) for the period 1950–2019.111The Conference Board Total Economy Database™ (April 2022) - Output, Labor and Labor Productivity, 1950-2022, downloaded from https://www.conference-board.org/data/economydatabase/total-economy-database-productivity on April 13, 2023. Wenz et al., (2023) provide subnational economic output data for over 1,661 regions across 83 countries, enabling panel and cross-sectional regression analyses that reduce coverage bias and increase the number of observations. Building on this approach, we spatially disaggregate TED’s country-level real GDP level into the real regional product level from 1950 to 2019. Using these data, we estimate panel fixed-effects models to remove persistent regional heterogeneity and long-term structural changes, and then relate the residual component of regional growth to climate variables. The detailed procedures for spatial disaggregation, density generation (on the support [0.105,0.092][-0.105,0.092], excluding a few extreme observations), and the panel specifications are provided in Appendix E.
@@
The literature notes that treating probability densities with support $a_{1},a_{2}]$ (without transformation) as Hilbert-space elements is inadvisable (e.g., Petersen and Müller,, 2016), since densities do not form a linear space; this issue is particularly pronounced for nonstationary density-valued time series (Seo and Beare,, 2019). To implement our framework, we apply the centered log-ratio (CLR) transformation of each density $g$, given by $g\mapsto\log g(u)-(a_{2}-a_{1})^{-1}\int_{a_{1}}^{a_{2}}\log g(s)ds$ which, under regularity conditions, maps (in a bijective manner) the density $g$ into the subspace of $L^{2}[a_{1},a_{2}]$ orthogonal to constant functions (Egozcue et al.,, 2006), enabling direct application of our methods. [^2] Since the CLR transformation involves $\log g(s)$, problems arise when $g(s)=0$. As is common in practice (e.g., Seo and Shang,, 2024), this is avoided by adding a small constant to $g(s)$. In our study, the densities are constructed on a restricted domain excluding a few extreme values, so this issue does not occur. The resulting CLR-transformed densities of GRP growth rates (resp. land temperature anomalies), generated from the raw data, are interpreted as measures of their true counterparts with measurement errors, and they are treated as functional data $y_{t}$ (resp. $\tilde{x}_{t}$) in (24). [^3] We assume that both time series include intercepts but no deterministic time trends, as in Chang et al., (2020).
@@ 5.2.2 Nonstationarity and testing procedure for $d_{N}$
We examine the nonstationarity of the CLR-transformed time series computed in the previous section and estimate the nonstationarity dimension $d_{N}$ of $\{x_{t}\}_{t\geq 1}$, an input to our inferential methods. We apply the variance-ratio testing procedure of Nielsen et al., (2023), shown to be robust to measurement errors (see Proposition C.1 and Remark C.1). The procedure determines $d_{N}$ by testing $H_{0}:d_{N}=d_{0}$ against $H_{1}:d_{N}<d_{0}$ sequentially for $d_{0}=5,\ldots,1$ until the null is not rejected for the first time (see Section C.2). The results, presented in Table 1, identify $d_{N}=2$ for the time series of $\tilde{x}_{t}$ (or $x_{t}$) at the standard 10% or 5% significance levels. For the proposed model to hold, the CLR-transformed densities of GRP growth rates must be nonstationary and have two or fewer stochastic trends (i.e., the nonstationarity dimension for $y_{t}$ must lie in $(0,2]$). Applying the same test, we obtain $p$-values for $d_{0}=5,\ldots,1$ of 0.3%, 1.3%, 6.3%, 62.6%, and 75.4%, strongly supporting the presence of two stochastic trends (since the procedure does not reject $H_{0}:d_{N}=2$ for the first time, concluding $d_{N}=2$ at a significance level $\alpha>6.3\%$).
@@ 5.2.3 Estimation results: economic impact of climate change
We present estimation results focusing on the economic impact of climate change. Our main interest is in the slope parameter $f$ in the model with an intercept (24). For the entire estimation procedure, we use the CLR-transformed time series, and the threshold $\alpha$ is determined as in our simulation experiments (Section B of the Appendix), yielding ${\mathsf{K}}=4$ in this empirical study. We use $d_{N}=2$, as estimated in Section 5.2.2.
@@
We compute the proposed estimator for $\kappa=1$ and $\kappa=0$ for comparison. The estimator $\widehat{f}_{\kappa}$ is an operator mapping one function to another. Although it can, in principle, be visualized (since $f$ is Hilbert–Schmidt in the present setup, the estimated Hilbert–Schmidt kernel can be plotted in three dimensions), such a plot is unlikely to yield meaningful insights for practitioners focused on the economic implications of climate-related scenarios or major events. Instead, we consider a functional change $\zeta$ in $x_{t}$, interpreted as a global warming shock to the world economy, and estimate its partial effect $f(\zeta)$ to quantify the economic damages resulting from the shock. The hypothetical global warming shock $\zeta$ can produce permanent effects, transitory effects, or both on regional economic growth. Permanent and transitory impacts are measured based on long-run (climate change) and short-run (inter-annual weather) variations of the functional changes, respectively. The estimated total-run response function thus illustrates how global warming collectively impacts the spatial distribution of regional growth rates. Note that while measurement error does not affect the consistency of the long-run response function for either $\kappa=0$ or $\kappa=1$, it does affect the consistency of the short-run response function (see Remark 4.6). Thus, setting a positive $\kappa$ is necessary for robust statistical inference on the total-run response function.
@@
To construct a representative global warming function, at first, we compute the mean difference between the first and second halves of the density estimates for the land temperature anomaly (hereafter, GW1). As shown in the left panel of Figure 2, global warming can be conceptualized in statistical terms as a probabilistic shift from negative to positive anomalies, capturing the long-run distributional change in the Earth’s land surface temperature over the past 70 years. Previous studies estimate the break date for the northern hemisphere temperature anomaly at 1985 in the NASA dataset and 1984 in the HadCRUT3 dataset (Estrada et al.,, 2013; Estrada and Perron,, 2019). Given the close similarity in statistical properties between the land temperature anomaly and the northern hemisphere series (Chang et al.,, 2020), we adopt 1985 as a credible break date marking the onset of global warming in GW1. Of course, practitioners consider an alternative conceptualization on the global warming. For example, one can define the global warming as the distributional shift over the first and last five years of the sample period (hereafter, GW2); see the right panel of Figure 2. In this setting, GW2 serves as a complementary measure, offering greater robustness to interannual variability and to uncertainties in the precise timing of the structural break. While Figure 2 shows the densities, we use the model with CLR-transformed densities due to mathematical issues noted in the literature (e.g., Petersen and Müller,, 2016; Seo and Beare,, 2019). Since the CLR map is bijective, we consider the CLR transformations of the densities in each panel of Figure 2 and define $\zeta$ as the difference between the CLR-transformed densities. This is treated as a global warming shock in the model.
@@
The estimate $\widehat{f}_{\kappa}(\zeta)$ captures the effect of the generated global warming shock on the CLR-transformed density of regional growth rates. Figure 3 presents the estimated total-run $(\widehat{f}_{\kappa}^{N}(\zeta))$ and short-run $(\widehat{f}_{\kappa}^{S}(\zeta))$ responses to the considered global warming shock. Since the regressor is likely contaminated by measurement error, statistical inference is conducted for the estimates with $\kappa=1$, using the theoretical results in Theorem 4.4 (and Corollary C.1 in the Appendix). Specifically, the local confidence interval is obtained by estimating the pointwise standard error from the residual covariance within a one-grid bandwidth neighborhood and scaling it by the normal critical value at each point (see Section 4.3 and (23)). The 95% confidence intervals for the locally averaged response functions indicate that, while the short-run effects are statistically insignificant, global warming has a significant total-run impact on regional economic growth (potentially due to the limited sample size and the slower convergence rate of $\widehat{f}_{\kappa}^{S}$ compared to $\widehat{f}_{\kappa}^{N}$).
@@
The downward slope of the total-run response function indicates that global warming reduces the share of regions with high temperature-related economic growth while increasing the share with lower growth. In other words, as land temperatures rise, the distribution of regional growth shifts toward weaker outcomes. The slope is generally steeper under GW2 than under GW1, indicating that the magnitude of the climate-induced shift in regional growth outcomes is more pronounced when global warming is defined by the first-versus-last 5-year contrast. When measurement error in the functional covariate is accounted for ($\kappa=1$), the slope of the total-run response function becomes steeper at the right tail compared to the case where the covariate is assumed free of measurement error ($\kappa=0$). This discrepancy likely reflects bias from measurement errors, implying that the magnitude of climate-related economic impacts is underestimated when such errors are ignored.
@@
From a practical perspective, it is more informative to visualize the distributional effect implied by $f(\zeta)$ in terms of changes in the probability density of GRP growth rates (noting that $f(\zeta)$ represents an effect on the CLR-transformed density). This is achieved by: (i) fixing a reference density and its CLR transform $y_{\text{ref}}$; and (ii) inverting the CLR-valued quantity $y_{\text{ref}}+\widehat{f}_{\kappa}(\zeta)$ back into the corresponding probability density, then comparing it with the reference density. For the inversion, the inverse CLR transformation, $g(s)\mapsto\exp(g(s))/\int_{a_{1}}^{a_{2}}\exp(g(u))du$, is applied (Egozcue et al.,, 2006). Furthermore, scaled global warming shocks $\zeta_{q}=q\zeta$ for $q\geq 0$ and their distributional effects are considered to examine how the reference density changes as the global warming shock intensifies or diminishes.
@@
The left and middle panels of Figure 4 show the result when the reference density is set to the average density of $y_{t}$ over the period 1951–1984, $q$ increases from 0 to $1.5$ and $\zeta$ is constructed from GW1 or GW2. As $q$ increases, both shocks shift the mass of the distribution leftward and modestly widen it, reflecting lower average growth rates and greater dispersion across regions. The right panel summarizes these changes in terms of the first two moments. The mean declines approximately linearly with $q$, while the variance increases at an accelerating rate. Across all scales, GW2 produces more pronounced changes than GW1 in both the mean and variance, indicating a stronger impact on the central tendency and dispersion of regional growth rates. Taken together, these results suggest that stronger global-warming shocks are associated with slower average growth and increased dispersion, demonstrating the usefulness of our approach as a practical tool for policymakers to evaluate the adverse economic impacts of climate change.
@@ 6 Concluding Remarks
This paper develops regression models for nonstationary and potentially error-contaminated functional time series and introduces a novel autocovariance-based inferential method. We believe the methodology is broadly applicable to problems involving nonstationary functional data. Not only to illustrate our approach, but also for its intrinsic importance, we apply our methodology to assess the economic impact of climate change. Our analysis provides empirical evidence that global warming has a negative effect on regional economic growth.
Supplementary Appendix

### v3

@@ 5.2 Empirical Applications: Economic Impact of Climate Change
We present an empirical study examining the economic impact of climate change using our proposed method. A large body of research in climate economics has shown that rising temperatures negatively affect economic growth (e.g., Dell et al.,, 2012; Burke et al.,, 2015; Newell et al.,, 2021; Cruz and Rossi-Hansberg,, 2023), and estimating these effects is of central importance because they inform the economic costs of climate change and policy responses. Motivated by this literature, we apply our proposed method to global climate and economic data to demonstrate its empirical usefulness in providing statistical evidence for this relationship along with broader estimates of the economic impact of climate change.
We consider appropriate transformations of the probability densities of gross regional product (GRP) growth rates ($y_{t}$), as a measure of regional economic activity, and land temperature anomalies ($x_{t}$), commonly used as indicators of climate change. Because these time series are expected to be nonstationary, contaminated by measurement errors, and to exhibit nonzero unconditional means, we employ model (24), with $u_{t}$ absorbing measurement errors in the dependent variable.
@@ 5.2.1 Raw data and functional data in analysis
We use non-infilled gridded land temperature anomaly data from a collaborative product of the Climatic Research Unit at the University of East Anglia, the Met Office Hadley Centre, and the National Centre for Atmospheric Science (CRUTEM.5.0.2.0, Osborn et al.,, 2021). We estimate spatially distributed temperature anomaly densities for 1951–2019 using a Gaussian kernel with Silverman’s bandwidth. To avoid COVID-19-related distortions in $y_{t}$, data from 2020 onward are excluded. At each time $t$, the distribution’s support is restricted to the range containing 99% of the total probability mass, [5.80,6.68][-5.80,6.68], thereby excluding outliers as in Chang et al., (2020).
For the GRP growth rates, we employ both the real GRP data of Wenz et al., (2023) for 1960–2019 and the real GDP in millions of 2021 international dollars, converted using Purchasing Power Parities, from the Conference Board Total Economy Database (TED) for 1950–2019.111The Conference Board Total Economy Database™ (April 2022) - Output, Labor and Labor Productivity, 1950-2022, downloaded from https://www.conference-board.org/data/economydatabase/total-economy-database-productivity on April 13, 2023. Wenz et al., (2023) provide subnational economic output data for over 1,661 regions across 83 countries, enabling panel and cross-sectional regression analyses that reduce coverage bias and increase the number of observations. Based on this approach, we spatially disaggregate TED’s country-level real GDP into the regional product levels from 1950 to 2019. Using these data, we estimate panel fixed-effects models to remove persistent regional heterogeneity and long-term structural changes, and then relate the residual component of regional growth to climate variables. Detailed procedures for spatial disaggregation, density generation (on support [0.105,0.092][-0.105,0.092], excluding extreme observations), and panel specifications are provided in Section E of the Supplementary Material.
@@
The literature notes that treating probability densities with support $a_{1},a_{2}]$ (without transformation) as standard Hilbert-valued elements is inadvisable (e.g., Petersen and Müller,, 2016), since densities do not form a linear space; this issue is particularly pronounced for nonstationary density-valued time series (Seo and Beare,, 2019). To implement our framework, we apply the centered log-ratio (CLR) transformation of each density $g$, given by $g\mapsto\log g(u)-(a_{2}-a_{1})^{-1}\int_{a_{1}}^{a_{2}}\log g(s)ds$ which, under regularity conditions, maps (in a bijective manner) the density $g$ into the subspace of $L^{2}[a_{1},a_{2}]$ orthogonal to constant functions (Egozcue et al.,, 2006), enabling direct application of our methods. [^2] Since the CLR transformation involves $\log g(s)$, problems arise when $g(s)=0$. As is common in practice (e.g., Seo and Shang,, 2024), this is avoided by adding a small constant to $g(s)$. In our study, the densities are constructed on a restricted domain excluding a few extreme values, so this issue does not occur. The resulting CLR-transformed densities of GRP growth rates (resp. land temperature anomalies), generated from the raw data, are interpreted as measures of their true counterparts with measurement errors, and they are treated as functional data $y_{t}$ (resp. $\tilde{x}_{t}$) in (24). [^3] We assume that both time series include intercepts but no deterministic time trends, as in Chang et al., (2020).
@@ 5.2.2 Nonstationarity and testing procedure for $d_{N}$
We examine the nonstationarity of the CLR-transformed time series computed in the previous section and estimate the nonstationarity dimension $d_{N}$ of $\{x_{t}\}_{t\geq 1}$, an input to our inferential methods. We apply the variance-ratio testing procedure of Nielsen et al., (2023), which is shown to be robust to measurement errors in Section C.2 of the Supplementary Material. The procedure determines $d_{N}$ by testing $H_{0}:d_{N}=d_{0}$ against $H_{1}:d_{N}<d_{0}$ sequentially for $d_{0}=d_{\max},\ldots,1$ until the null is not rejected for the first time, where $d_{\max}$ is a prespecified upper bound for $d_{N}$. The results with $d_{\max}=5$, presented in Table 1, identify $d_{N}=2$ at the standard 10% or 5% significance level. This testing procedure can also be useful for a different purpose in our empirical analysis. Our model requires that the stochastic trends of $y_{t}$ are explained by those of $x_{t}$ ($f^{N}$ captures this). A straightforward extension of this testing procedure can be used as a diagnostic check to see whether this is the case in practice, and the testing results are supportive of our empirical analysis (see Remark C.2 of the Supplementary Material for detailed discussion and results).
@@ 5.2.3 Estimation results: economic impact of climate change
We present estimation results on the economic impact of climate change. Our main interest is in the slope parameter $f$ in the model with an intercept (24). The estimation uses the CLR-transformed time series, with the threshold $\alpha$ set as in our simulation experiments (Section B.2 of the Supplementary Material), yielding ${\mathsf{K}}=4$ in this study. We use $d_{N}=2$, as estimated in Section 5.2.2. Since $f^{N}$ can be viewed as capturing a stable long-run relationship between persistent stochastic trends, it is called the long-run response function, while $f^{S}$ is called the short-run response function, as in the literature (e.g., Meierrieks and Stadelmann,, 2023); accordingly, $f$ is referred to as the total-run response function.
@@
We compute the estimators for $\kappa=1$ and $\kappa=0$ for comparison. The estimator $\widehat{f}_{\kappa}$ is an operator that maps a function to another. Although it can be visualized (since $f$ is Hilbert–Schmidt in the present setup, the estimated Hilbert–Schmidt kernel can be plotted in three dimensions), such a plot is unlikely to yield meaningful insights for practitioners concerned with the economic implications of climate-related scenarios or major events. Instead, we consider a functional change $\zeta$ in $x_{t}$, interpreted as a global warming shock to the world economy, and estimate its partial effect $f(\zeta)$ to quantify the economic damages resulting from the shock. The hypothetical global warming shock $\zeta$ can produce permanent effects, transitory effects, or both on regional economic growth. Permanent and transitory impacts are measured based on long-run (climate change) and short-run (interannual weather) variations of the functional changes, respectively. The estimated total-run response function thus illustrates how global warming collectively impacts the spatial distribution of regional growth rates. Note that while measurement errors do not affect the consistency of the long-run response function estimator $\widehat{f}_{\kappa}^{N}$ for either $\kappa=0$ or $\kappa=1$, they do affect the consistency of the short-run response function estimator $\widehat{f}_{\kappa}^{S}$ (see Remark 4.6). Thus, setting a positive $\kappa$ is necessary for robust statistical inference on the total-run response function.
@@
To construct a representative global warming function, at first, we compute the mean difference between the first and second halves of the density estimates for the land temperature anomaly (hereafter, GW1). As shown in the left panel of Figure 2, global warming can be conceptualized in statistical terms as a probabilistic shift from negative to positive anomalies, capturing the long-run distributional change in the Earth’s land surface temperature over the past 70 years. Previous studies estimate the break date for the northern hemisphere temperature anomaly at 1985 in the NASA dataset and 1984 in the HadCRUT3 dataset (Estrada et al.,, 2013; Estrada and Perron,, 2019). Given the close similarity in statistical properties between the land temperature anomaly and the northern hemisphere series (Chang et al.,, 2020), we adopt 1985 as a credible break date marking the onset of global warming in GW1. Of course, practitioners consider an alternative conceptualization of global warming. For example, one can define global warming as the distributional shift over the first and last five years of the sample period (hereafter, GW2); see the right panel of Figure 2. In this setting, GW2 serves as a complementary measure, offering greater robustness to interannual variability and to uncertainties in the precise timing of the structural break. While Figure 2 shows the probability densities, we use the model with CLR-transformed densities due to mathematical issues noted in the literature (e.g., Petersen and Müller,, 2016; Seo and Beare,, 2019). Since the CLR map is bijective, we consider the CLR transformations of the densities in each panel of Figure 2 and define $\zeta$ as the difference between the CLR-transformed densities. This is treated as a global warming shock in the model.
@@
The estimate $\widehat{f}_{\kappa}(\zeta)$ captures the effect of the generated global warming shock on the CLR-transformed density of regional growth rates. Figure 3 presents the estimated total-run $(\widehat{f}_{\kappa}(\zeta))$ and short-run $(\widehat{f}_{\kappa}^{S}(\zeta))$ responses to the considered global warming shock. Since the regressor is likely contaminated by measurement errors, statistical inference is conducted for the estimates with $\kappa=1$, using the theoretical results in Theorem 4.4 (and Corollary C.1 in the Supplementary Material). Specifically, the local confidence interval is obtained by estimating the pointwise standard error from the residual covariance within a one-grid bandwidth neighborhood and scaling it by the normal critical value at each point (see Section 4.3 and (23)). The 95% confidence intervals for the locally averaged response functions indicate that, while the short-run effects are statistically insignificant, global warming has a significant total-run impact on regional economic growth (potentially due to the limited sample size and the slower convergence rate of $\widehat{f}_{\kappa}^{S}$ compared to $\widehat{f}_{\kappa}^{N}$).
@@
The downward slope of $\widehat{f}_{\kappa}(\zeta)$ indicates that global warming reduces the share of regions with high temperature-related economic growth while increasing the share with lower growth. In other words, as land temperatures rise, the distribution of regional growth shifts toward weaker outcomes. The slope is generally steeper under GW2 than under GW1, indicating that the magnitude of the climate-induced shift in regional growth outcomes is more pronounced when global warming is defined by the first-versus-last 5-year contrast. When measurement errors in the functional data are accounted for (with $\kappa=1$), the slope of $\widehat{f}_{\kappa}(\zeta)$ becomes steeper at the right tail compared to the error-free case (with $\kappa=0$). This discrepancy likely reflects bias from measurement errors, implying that the magnitude of climate-related economic impacts is underestimated when such errors are ignored.
@@
From a practical perspective, it is more informative to visualize the distributional effect implied by $f(\zeta)$ in terms of changes in the probability density of GRP growth rates (noting that $f(\zeta)$ represents an effect on the CLR-transformed density). This is achieved by (i) fixing a reference density and its CLR transformation $y_{\text{ref}}$, and (ii) inverting the CLR-valued quantity $y_{\text{ref}}+\widehat{f}_{\kappa}(\zeta)$ back into the corresponding probability density, then comparing it with the reference density. For the inversion, the inverse CLR transformation $g(s)\mapsto\exp(g(s))/\int_{a_{1}}^{a_{2}}\exp(g(u))du$ is applied (Egozcue et al.,, 2006). Moreover, scaled global warming shocks $\zeta_{q}=q\zeta$ for $q\geq 0$ and their distributional effects are considered to examine how the reference density changes as the global warming shock intensifies or diminishes.
@@
The left and middle panels of Figure 4 show the result when the reference density is set to the average density of $y_{t}$ over the period 1951–1984, $q$ increases from 0 to $1.5$, and $\zeta$ is constructed from GW1 or GW2. As $q$ increases, both shocks shift the mass of the distribution leftward and modestly widen it, reflecting lower average growth rates and greater dispersion across regions. The right panel summarizes these changes in terms of the first two moments. The mean declines approximately linearly with $q$, while the variance increases at an accelerating rate. Across all scales, GW2 produces more pronounced changes than GW1 in both the mean and variance, indicating a stronger impact on the central tendency and dispersion of regional growth rates. Taken together, these results suggest that stronger global-warming shocks are associated with slower average growth and increased dispersion, demonstrating the usefulness of our approach as a practical tool for policymakers to evaluate the adverse economic impacts of climate change.
@@ 6 Concluding Remarks
This paper develops regression models for nonstationary and potentially error-contaminated functional time series and introduces a novel autocovariance-based inferential method. The methodology is broadly applicable to problems involving nonstationary functional data. Not only to illustrate our approach, but also for its intrinsic importance, we apply our methodology to assess the economic impact of climate change. Our analysis provides empirical evidence that global warming has a negative effect on regional economic growth.
Appendix